If an ideal gas with an initial volume of 25 mL and pressure of 0.99 atm increases its volume to 45 mL, what is its new pressure? and which gas law did u use to solve the question above?


1
Expert's answer
2021-03-08T07:09:30-0500

Using Boyle law

P1/P2= V2/V1

0.99/p2 = 45/25

p2 =0.55atm